OpenStudy (anonymous):

what equation would represent a line that is paralel to the points (-5.6) (-3,-4)

OpenStudy (anonymous):

is this a multiple choice question? There are many many many different answers

OpenStudy (anonymous):

literally infinite amount of answers

OpenStudy (anonymous):

yes, sorry guys, y=-5x-22 y=1/5x -22 y=-1/5x -22 y=5x-22

OpenStudy (anonymous):

the answer is y = -5x - 22

OpenStudy (anonymous):

oh thanks!